(-2,-77, find -10r For the vector r:
Answer #1


Answer veetor y = <-27-7> -2-7j -108 - -] -8=-(-20 - 7)) -8= 21+ 7j - 107 10 r2 i +7 1 ) -108 - 201 + 7oj لوک -107 < 20, 70 >

Please upvote this answer

If you have any doubts then comment below